Testing on May 29, 2014

Plan:

  • Try out new bof file for modes 10-19 and see if the image is gone
  • Double-check that mode 20-29 are okay if mode 10-19 are okay

Info:

  • TGBT13B_502_64
  • Amanda in control room

Log:

  • scan 1: fsw_issue noswitch_mode10. Trying one of the problem modes. Yep. Still see the image
  • scan 2: fsw_issue_noswitch_nocal_mode20. Trying one of the good modes. Still looks good.
  • Let Paul have it for his tests.
Topic revision: r2 - 2015-01-27, RichardPrestage
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback